Write down all derangements of the set \left\{ a,b,c,d \right\} and show that the number of derangements is the same as predicted by the recurrence D(n) = (n - 1)(D(n - 2) + D(n - 1)) with initial values D(1) = 0 and D(2) = 1. Hint: a derangement is a permutation of an ordered set where no element is in the same place as before. Example: \left\{ b,a,d,c \right\} is a derangement of \left\{ a,b,c,d \right\} because all of the letters positions have changed.

1
Expert's answer
2022-03-01T04:07:02-0500

There are nine derangements for the set "\{a,b,c,d\}". The derangements are

"\left\{\{b,a,d,c\}, \{b,c,d,a\}, \{b,d,a,c\}, \{c,a,d,b\}, \{c,d,a,b\}, \{c,d,b,a\}, \{d,a,b,c\},\right.\\ \left. \{d,c,a,b\}, \{d,c,b,a\}\right\}."


Let "D(n)" denote the number of derangements of n distinct objects.

For each derangement, let 1 be placed in position "i" for "2\le i\le n". We have two cases to consider.


Case 1: "i" is in position 1

Then there are "D(n - 2)" ways to derange the remaining "n-2" integers. Since there are "n - 1" choices for "i" we have "(n-1)D(n-2)" derangements.


Case 2: "i" is not in position 1

Then there are "D(n-1)" ways to derange the "n - 1" integers by considering 1 as new natural position for "i". Again with "n -1" choices for "i", we have "(n-1)D(n-1)" derangements.


Hence, "D(n) = (n-1)(D(n-2)+D(n-1))" with "D(1) =0~ \& ~D(2) = 1", since there are no derangements possible for 1 object and exactly one derangement for 2 objects.


Now, "D(4) = (4-1)(D(2)+D(3)) = 3(D(2) + 2(D(1)+D(2))) = 3(1+2)=9".

Thus, the number of derangements is the same as predicted by the recurrence relation.


Note: Book for reference is "DISCRETE AND COMBINATORIAL MATHEMATICS - An Applied Introduction - FIFTH EDITION by RALPH P. GRIMALDI".